OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10, 5/3, -5/18,

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

might help to see what the tops look like when they have like denominatos

OpenStudy (anonymous):

so all have denominator 18?

OpenStudy (amistre64):

1080, -180, 30, -5 ------------------ 18 well, the pattern seems to be wobbling back and forth between signs and getting smaller and smaller

OpenStudy (anonymous):

ahhh I see so tht would be diverging?

OpenStudy (amistre64):

and if miscalculated some of them :)

OpenStudy (anonymous):

huh? :D

OpenStudy (amistre64):

since they are bobbling back and forth, lets just absolute value it 1080, 180, 90, 5 --------------- 18 can we define a expression for the top sequence?

OpenStudy (amistre64):

since this is clearly not arithmetic ....dividing by 5 we get {1080, 180, 90, 5}/5 = { 216, 36, 18, 1}

OpenStudy (amistre64):

divide by 6: 36, 6 ,3 ,1/6 divide by 3: 12, 2 ,1 ,1/18 divide by 2: 6, 1 ,1/2 ,1/36 divide by 6: 1, 1/6 ,1/12 ,1/236 in some fashion or another the top is approaching 0

OpenStudy (amistre64):

can you find a closed form to operate with?

OpenStudy (amistre64):

60/6 = 10 10/6 = 5/3 (5/3)/6 = 5/18 its a geometric series

OpenStudy (amistre64):

\[\sum_0~\frac{(-1)^n~60}{6^n}\]maybe

OpenStudy (anonymous):

wait I am so lost!:(

OpenStudy (amistre64):

this might be helpful http://www.wolframalpha.com/input/?i=table+%28-1%29%5En*60%2F6%5En%2C+n%3D0+to+10

OpenStudy (anonymous):

so it would be converging at 0?

OpenStudy (amistre64):

the limit of the sequence is: \[\lim_{n\to inf}~\frac{60~(-1)^{n+1}}{6^{n+1}}*\frac{6^{n}}{60~(-1)^{n}}\] \[\lim_{n\to inf}~\frac{(-1)^{n+1-n}}{6^{n+1-n}}\] \[\lim_{n\to inf}~\frac{(-1)}{6}\] is that correct?

OpenStudy (anonymous):

well that's not an option the answers are set up as Diverges Converges; 11100 Converges; 72 Converges; 0

OpenStudy (amistre64):

lol, my latex coding is off .... the limit of the sequence goes to zero which indicates that the series has hope for convergence

OpenStudy (anonymous):

Excellent! That's what I was thinking when I saw wolfram! :)

OpenStudy (amistre64):

im reading something into it thats not there ... i was trying for convergent series for some odd reason :/

OpenStudy (anonymous):

I see, :you were great either way :') thanks so much!!!

OpenStudy (amistre64):

youre welcome

OpenStudy (nikvist):

sequence diverges, because is not monotone

OpenStudy (anonymous):

what??

OpenStudy (nikvist):

sequence converges, when is monotone and has limit

OpenStudy (anonymous):

so it can't be converges at 0??

OpenStudy (amistre64):

\[a_n = 60(\frac{-1}{6})^n\] \[\lim_{n\to \inf}|a_n| = 0\]

OpenStudy (anonymous):

Yeah.. that seems accurate!

OpenStudy (amistre64):

since |-1/6| < 1 .... it goes to zero 60*0 = 0

OpenStudy (anonymous):

Yep! it'sdefinitely correct, thanks loads! @amistre64

OpenStudy (amistre64):

youre welcome ... again :)